OpenStudy (anonymous):

The areas of two similar triangles are 16cm^2 and 25 cm^2. One of the sides of the second triangle is 8cm. What is the length of the corresponding side of the other triangle?

OpenStudy (anonymous):

@Luis_Rivera You did it wrong..

OpenStudy (anonymous):

@kropot72 Can you help me

OpenStudy (anonymous):

4/5=8/x x=10

OpenStudy (anonymous):

but 4/5 is the ratio of similitude

OpenStudy (aravindg):